# A Comparative Study of Conventional and Tripolar EEG for High-Performance Reach-to-Grasp BCI Systems

### v1

@@ 4 Discussion
This study corroborates the enhanced proficiency of tEEG over conventional EEG in grasping movement decoding, which is crucial for BCI systems. Aligning with prior research [31, 32], tEEG showed superior spatial resolution and signal-to-noise ratio, with reduced channel crosstalk in functional connectivity and ERP analyses. Further, the ERP data from both conventional EEG and tEEG reveal a distinct neural response to the movement onset across all conditions. However, tEEG seems to offer a more detailed picture, potentially due to its enhanced spatial resolution. The multiplicity of traces in tEEG, especially for dynamic conditions like Power and Precision Grasp, underscores its capability to capture a broader spectrum of neural responses. In the No Movement condition, the disparity between the two methodologies is most evident. While conventional EEG depicts a steady state, tEEG offers a mosaic of neural activities, suggesting that even in a static condition, the brain is far from inactive.
Uniquely, wavelet time-frequency maps indicated that tEEG captured more pronounced brain activity post-movement. This facilitated the improved accuracy of motor task classification across machine learning models, as demonstrated in the performance enhancements.
The enhanced decoding of grasp movements via tEEG in this study not only aligns with previous research but also surpasses it. tEEG’s benefits in binary classification tasks were particularly significant, with Random Forest and XGBoost showing major accuracy gains. Although SVM performed robustly with both EEG types,tEEG’s enhancements were less pronounced, indicating a need for further study into feature selection for SVM. Importantly, our work achieves up to 91.53% accuracy for No-movement vs. Power Grasp, 90.71% for No-movement vs. Precision Grasp, and 87.88% for multiclass tasks, demonstrating comparable or superior performance to existing research [26, 27, 28], even with cost-effective EEG headset. The tEEG’s advantage is most evident in multiclass classification, where it substantially outperforms conventional EEG. Specifically, improvements of 25.45% for RF, 23.87% for XGB, and 22.62% for LDA models highlight tEEG’s capability in handling complex neural pattern distinctions crucial for advanced BCI applications, supporting the necessity for high-fidelity signal capture [50, 51].
@@ 5 Conclusion
In conclusion, this study affirms tEEG’s superior capability in decoding grasp movements for BCI applications. Leveraging its higher spatial resolution and signal-to-noise ratio, tEEG outperformed conventional EEG in binary and multiclass classification tasks, supported by notable accuracy gains with machine learning models. The advancements in tEEG technology showcase its potential to enhance BCI functionality, even with cost-efficient setups. Future research should delve into optimizing tEEG for individual neural patterns and exploring its efficacy across various BCI paradigms requiring intricate motor control.
6 Data availability statement
The data collected and analyzed in this study are not openly accessible due to legal and ethical constraints. However, they can be obtained from the corresponding author upon a reasonable request.
@@ 7 Acknowledgments
This Research was partially supported by the Rhode Island INBRE program from the National Institute of General Medical Sciences of the National Institutes of Health under grant number P20GM103430 and the National Science Foundation under award ID 2245558. Preliminary results of this study were presented at the Society for Neuroscience 2023 Meeting [52] (November 2023, Washington DC, USA).
8 Conflict of interest
The authors declare that there are no relevant conflicts of interest associated with this paper.

### v2

@@ 4 Discussion
This study corroborates the enhanced proficiency of tEEG over conventional EEG in grasping movement decoding, which is crucial for BCI systems. Aligning with prior research [aghaei2020tripolar], tEEG showed an improved signal-to-noise ratio across various channels. This indicates tEEG’s superiority in producing clearer, more reliable EEG signals. Specifically, as it is illustrated in Figure 3, tEEG showed higher median SNRs, especially notable at channel P3, and demonstrated less variability among subjects, suggesting more consistent signal quality. Improved signal-to-noise ratio (SNR) in electroencephalography (EEG) signals can significantly enhance the performance of brain-computer interfaces (BCIs) by increasing the separability of neural patterns associated with different mental tasks or intentions [gu2021eeg]. Higher SNR leads to better feature extraction and classification accuracy, resulting in more reliable and efficient BCI systems. [assi201733] Further, the results in Figure 4 indicate that the tEEG technology exhibits significantly lower functional connectivity between EEG channels compared to conventional EEG, and this finding is in line with prior research [koka2007improvement]. The heatmaps show substantially weaker correlations among channel pairs for tEEG, indicating higher spatial resolution and improved ability to localize and distinguish neural activities originating from different brain regions. This lower functional connectivity and higher spatial resolution afforded by tEEG can potentially enhance the performance of brain-computer interface (BCI) systems by providing more precise and localized information about task-related neural activities.
In addition, the observed differences in the time-frequency representations between EEG and tEEG, shown in Figure 5, underscore the potential of tEEG to provide significantly improved temporal and spectral resolution. This advancement could facilitate more precise identification of distinct neural patterns associated with various motor tasks, thereby contributing to the improved accuracy of motor task classification across diverse machine learning models employed in this study. Notably, the advantages of tEEG were particularly pronounced in binary classification tasks, where Random Forest and XGBoost classifiers exhibited substantial gains in accuracy when leveraging tEEG signals compared to conventional EEG. Although the Support Vector Machine (SVM) classifier demonstrated robust performance with both EEG modalities, the enhancements afforded by tEEG were less pronounced, suggesting a need for further investigation into optimal feature selection strategies tailored for SVM classifiers when utilizing tEEG data. Importantly, our work achieves up to 91.43% accuracy for no-movement vs. power grasp, 90.71% for no-movement vs. precision grasp, 87.88% for power vs. precision grasp, and 75.97% for multiclass tasks demonstrating comparable or superior performance to existing research [iturrate2018human, schwarz2017decoding, xu2021decoding], even with cost-effective EEG headset. The tEEG’s advantage is most evident in multiclass classification, illustrated in Figure 6, where it significantly outperforms conventional EEG. Specifically, improvements of 25.45% for RF, 23.87% for XGB, and 22.62% for LDA models highlight tEEG’s capability in handling complex neural pattern distinctions crucial for advanced BCI applications, supporting the necessity for high-fidelity signal capture [saha2021progress, roman2012eeg]. These results underscore the robustness of the features extracted using wavelet-based techniques and tEEG signals, which consistently enable superior performance across various machine learning algorithms. This demonstrates the effectiveness of combining tEEG technology with advanced signal processing techniques to enhance the decoding of complex neural activities in grasp-related tasks.
The findings from this study underscore the significant potential of tEEG technology to improve BCI systems. By offering enhanced signal quality and reliability, higher spatial and temporal resolutions, and improved discrimination of neural activities, tEEG enables more accurate and nuanced control within BCI applications. Furthermore, the improved performance in motor task classification, as demonstrated across various machine learning models, suggests that tEEG could facilitate the development of BCIs with broader applications, from medical rehabilitation to control of external devices. Essentially, these findings highlight tEEG’s role in pushing the boundaries of what’s achievable with current BCI technology, offering a pathway to more sophisticated and accessible interfaces that could significantly enhance user interaction with technology.
5 Limitations and Future Works
While our findings are promising, it is important to highlight several considerations. The better signal‑to‑noise ratio and sharper spatial resolution of tripolar EEG (tEEG) were already shown in other studies [besio2006tri], so their presence here is not surprising. We further contributed by applying wavelet-based time-frequency analysis to show how tEEG enhances the spectral and temporal characterization of grasp-related neural dynamics. Additionally, we performed the first direct test of how much these hardware gains translate into higher decoding accuracy for grasp-related brain-computer-interface (BCI) tasks.
However, several limitations should be acknowledged. First, our study involved only ten healthy adults between 22 and 35 years old, a relatively small sample that does not capture the broader variability found in clinical or aging populations. Second, data were collected from only four scalp locations at a sampling rate of 250 Hz, limiting our ability to detect high-gamma activities and potentially missing important signals from other motor or parietal regions. Third, the task paradigm included only three categories—power grasp, precision grasp, and rest, whereas real-world grasping behaviors involve a wider range of forces, intentions, and movement complexities.
To address these limitations, future work will pursue several directions. First, we will conduct a larger-scale study that includes a more diverse participant pool spanning different age groups and clinical populations. Second, we aim to expand the tEEG montage beyond four electrodes and increase the sampling rate to capture higher-frequency components, such as high-gamma activity, and to better characterize signals from additional motor and parietal regions involved in grasp planning and execution. Third, we plan to extend the task design to include a wider range of grasp types and force levels (e.g., soft, medium, hard), as well as intentional “null” or “no-go” grasps. These additions will enable the development of a more fine-grained tEEG-based decoding pipeline, better aligned with the complexity of real-world grasping behaviors.
@@ 6 Conclusion
This work delivers systematic evidence that the well‑known hardware advantages of tripolar EEG translate into meaningful performance gains for practical brain‑computer interfaces. Compared with conventional disc electrodes, tEEG raised signal‑to‑noise ratios, lowered spurious channel‑to‑channel correlations, and, most importantly, boosted classification accuracy by 12–25 percentage points across binary and multiclass grasp‑decoding tasks. These improvements were achieved with just four scalp sites and lightweight machine‑learning models, suggesting that even modest tEEG headsets can outperform far denser conventional arrays.
Beyond technical merit, this study provides a practical roadmap for scaling tEEG from laboratory prototypes to everyday devices. By quantifying SNR gains, spatial‑resolution benefits, and decoder accuracy in a single framework, we furnish design targets for hardware engineers and algorithm developers alike. As future work extends tEEG testing to larger and clinical cohorts, incorporates richer grasp taxonomies, and validates real‑time performance, tripolar electrodes are poised to shift non‑invasive BCIs from proof‑of‑concept toward routine, user‑friendly solutions that bridge the gap between neural intention and the physical world.
@@ 7 Acknowledgments
The data can be obtained from the corresponding author upon a reasonable request. Preliminary results of this study were presented at the Society for Neuroscience 2023 Meeting [rabiee2023decoding] (November 2023, Washington DC, USA).